OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 4x - y = 4 x + 4y = 18 Choose one answer. A. (2, 4) B. (4, 2) C. (1, 4) D. (4, 1)

OpenStudy (anonymous):

Ok. X+4y =18; x = 18-4Y 4(18-4Y) - Y = 4 72 - 16Y - Y = 4 72 - 17Y = 4 72 - 4 = 17Y 68 = 17Y; Y= 4 The correct answer HAS to be A.